OpenStudy (anonymous):

If $240 is invested at an interest rate of 9% per year and is compounded monthly, how much will the investment be worth in 14 years?

OpenStudy (anonymous):

\(\LARGE\color{black}{ \bf A=240(1+\frac{0.09}{12})^{(12 \times 14)} }\)